### v2

# QWENDY: Gene Regulatory Network Inference Enhanced by Large Language Model and Transformer

Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. Based on a linear gene expression model, it solves the transformation for the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. Then we enhance QWENDY by the transformer neural networks and large language models to obtain two variants: TEQWENDY and LEQWENDY. We test the performance of these methods on two experimental data sets and two synthetic data sets. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets.
1 Introduction
One gene can activate or inhibit the expression of another gene. Genes and such regulation relations form a gene regulatory network (GRN). For $n$ genes, the corresponding GRN is commonly expressed as an $n\times n$ matrix $A$, where $A_{ij}>0/=0/<0$ means gene $i$ has positive/no/negative regulation effect on gene $j$, and its absolute value represents the regulation strength. If the regulation relations of concerned genes are known, one can understand the corresponding biological process or even control it with gene perturbation. Therefore, knowledge of GRNs can be useful in developmental biology [1, 2, 3], and even in the study of macroscopic behavior [4, 5, 6]. Since it is difficult to determine the GRN directly, the common practice is to infer the GRN from gene expression data.
With new measurement techniques, such as scRNA-seq, one can measure the expression levels (mRNA counts) of different genes for a single cell. Since gene expression at the single cell level is random, this measurement can be repeated for different cells, and we can use those samples to obtain the probability distribution of different genes at this time point. Since the measurement kills cells, one cell can only be measured once. Therefore, we cannot obtain the joint probability distribution of gene expression levels at different time points. When we measure at multiple time points, we can only obtain a marginal probability distribution of different genes for each time point.
If the gene expression is at stationary, measurement at multiple time points will produce the same distribution. If the gene expression is away from stationary, such as after adding drugs or during development, we can measure the gene expression at multiple time points and obtain several different probability distributions. For single-cell expression data of $n$ genes at one time point, we can calculate the mean value of each gene ($n$ independent values) and the $n\times n$ covariance matrix ($n(n+1)/2$ independent values, since the covariance matrix is symmetric), while higher-order statistics are not numerically stable due to limited cell number. Therefore, we have $n+n(n+1)/2$ independent known values, much smaller than what is needed to fully determine the $n\times n$ GRN [7]. Therefore, we prefer the non-stationary distributions measured at multiple time points, since they contain enough information to infer the GRN.
For such single-cell gene expression data measured at multiple time points, although it is the most informative data type under current technology, there are only a few GRN methods developed specifically for this data type [8]. We have developed WENDY method to infer the GRN with this data type [9]. It needs measurement at two time points, and solves the transformation between the covariance matrices of two probability distributions, where the transformation is determined by the GRN. However, it needs to solve a non-convex optimization problem, which has infinitely many solutions, and WENDY will output one solution, determined by the numerical solver.
In this paper, we present an improved version of WENDY that needs data measured at four time points, and it is also based on solving the transformation between covariance matrices. This new method is named QWENDY, where Q stands for quadruple. With the help of data from more time points, QWENDY can uniquely determine the GRN. Besides, QWENDY does not need to conduct non-convex optimizations, but just matrix decompositions.
WENDY method is derived from a simplified gene expression model, and the performance of WENDY is not satisfactory when the model does not fit with experiments. Instead of developing a more complicated gene expression model, we developed TRENDY method [10], which trains a transformer model, a deep learning architecture that can be applied to various fields [11], to transform the input data to better fit the gene expression model. Then the transformed data will produce more accurate GRNs by WENDY. Such more accurate GRNs will be further enhanced by another transformer model.
We use the idea of TRENDY to enhance QWENDY. The enhanced version of QWENDY that trains two new transformer models is named TEQWENDY. Besides, we also present another approach that fine-tunes a large language model (LLM) to replace the transformer model, since it has a large pre-trained transformer section. This LLM-enhanced version is named LEQWENDY. The training of TEQWENDY and LEQWENDY uses synthetic data.
We test the performance of QWENDY, TEQWENDY, and LEQWENDY on two experimental data sets and two synthetic data sets. The TRENDY paper [10] has tested 16 methods on the same data sets, meaning that we can compare them with our three new methods. In all 19 GRN inference methods, TEQWENDY has the best overall performance. If we restrict ourselves to two experimental data sets, QWENDY ranks first. Therefore, we recommend applying QWENDY or TEQWENDY.
Section 2 reviews other GRN inference methods, especially those related to LLMs. Section 3 introduces QWENDY method. We develop TEQWENDY and LEQWENDY methods in Section 4. Section 5 tests the performance of different methods on four data sets. We finish with some discussions in Section 6.
2 Literature review
There have been numerous non-deep learning GRN inference methods, and they can be roughly classified as information-based and model-based. Information-based methods [12, 13, 8] do not rely on models for gene expression, but treat GRN inference as a feature selection problem: for a target gene, select genes that can be used to predict the level of the target gene. A common obstacle for information-based methods is to distinguish between direct and indirect regulations.
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [14] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [15, 16] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [17] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, where each cell is measured only once, we only know one model-based method, WENDY.
Deep learning-based GRN inference methods [18, 19, 20, 21, 22] generally use neural networks as black boxes, without integrating them with gene expression models. Therefore, lacking of interpretability is a common problem.
Similar to TRENDY, there are some approaches to enhance GRNs inferred by other known methods[23, 24, 25, 26, 27], but they generally require extra knowledge of transcription factors or cannot fully determine the GRN.
Since WENDY paper and TRENDY paper contain more detailed summaries for traditional GRN inference methods and deep learning-based GRN inference methods, we present a review for GRN inference methods related to LLMs.
Some researchers directly used LLMs as an oracle machine to generate answers, without further training the model. Azam et al. [28] asked different GRNs whether one gene regulates another gene. Afonja et al. [29] provided GPT-4 with potential transcription factors and asked it to generate a GRN. Wu et al. [30] provided GPT-4 with related papers and asked it to summarize regulations and form a GRN.
Some researchers worked on pre-trained LLMs and fine-tuned them with new data. Weng et al. [31] trained GPT-3.5 with related papers to obtain GenePT, a new LLM that can provide a high-dimensional embedding (a vector of real numbers) of each gene. Then this embedding was used to train a neural network to output the GRN. Yang et al. [32] trained BERT with scRNAseq data to obtain scBERT, which can also provide gene embedding. Kommu et al. [33] used the embedding from scBERT to infer the GRN.
Some researchers trained new models from scratch. Cui et al. [34] trained a new model, scGPT, with expression data from many cells. The structure of scGPT is similar to other LLMs, although with a smaller size. This model can provide gene embedding, which was used to infer GRN.
For general applications of LLMs in bioinformatics, readers may refer to two reviews [35, 36].
3 QWENDY method
3.1 Setup
At four time points $T=0$, $T=t$, $T=2t$, $T=3t$, we measure the single-cell gene expression levels. For the expression levels of $n$ genes from $m$ cells at each time point, we treat them as $m$ samples of an $n$-dimensional probability distribution. Then we first calculate the average level of each gene over $m$ cells, and use graphical lasso to calculate the $n\times n$ covariance matrix for different genes. The $1\times n$ expected levels at four time points are denoted as $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$. The covariance matrices at four time points are denoted as $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$. Notice that the process does not start from stationary, so that the probability distributions of $n$ genes are different for different time points, and these $\mathbf{x}_{i}$ and $K_{i}$ are not equal.
In the WENDY paper, the relationship for $K_{0},K_{1}$ and the GRN $A$ is derived after some approximations:
$K_{1}=(I+tA^{\mathrm{T}})K_{0}(I+tA),$
where $I$ is the $n\times n$ identity matrix. Define $B=I+tA$, we have
$K_{1}=B^{\mathrm{T}}K_{0}B.$ (1)
From this equation, WENDY directly solves $B$ (and thus $A$) by minimizing $||K_{1}-B^{\mathrm{T}}K_{0}B||_{\mathrm{F}}^{2}$, where $\mathrm{F}$ is the Frobenius norm. This problem is non-convex and has infinitely many solutions. WENDY outputs one solution of them.
For $K_{2},K_{3}$, similarly, we also have approximated equations
$K_{2}=B^{\mathrm{T}}K_{1}B,$ (2)
$K_{3}=B^{\mathrm{T}}K_{2}B.$ (3)
In this paper, we study whether we can better solve $B$ with data from more time points, especially $K_{2},K_{3}$. Since $B^{\mathrm{T}}KB=(-B)^{\mathrm{T}}K(-B)$, we cannot distinguish between $B$ and $-B$ from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$.
To solve this problem, we use $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$. In the WENDY paper, it is derived that
$\mathbf{x}_{1}=\mathbf{x}_{0}B+t\mathbf{c},$ (4)
where $\mathbf{c}$ is an unknown vector. Similarly,
$\mathbf{x}_{2}=\mathbf{x}_{1}B+t\mathbf{c},$ (5)
$\mathbf{x}_{3}=\mathbf{x}_{2}B+t\mathbf{c}.$ (6)
Since Eqs. 16 are approximated, there are two problems: (1) if these equations hold accurately, can we solve $B$; (2) if these equations do not quite hold, can we find $B$ to minimize the error. We will present the QWENDY method that provides positive answers to both problems.
For the first problem, Theorem 1 proves that given $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$, $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ that satisfy Eqs. 16 for some $B_{0}$, QWENDY can solve $B_{0}$ uniquely.
For the second problem, there are different interpretations.
One interpretation is to minimize
$||K_{1}-B^{T}K_{0}B||_{F}^{2}+||K_{2}-B^{T}K_{1}B||_{F}^{2}+||K_{3}-B^{T}K_{2}B||_{F}^{2}$
for any $B$. Unfortunately, this optimization problem is non-convex, making it difficult to solve.
Notice that for the gene expression data after interpretation, $K_{0}$ is measured earlier than $K_{3}$, making it farther from stationary, and more informative. Therefore, we want to emphasize more on $K_{0}$, $K_{1}$ than $K_{2}$, $K_{3}$. Our goal is to first find $B$ that minimizes $K_{1}-B^{\mathrm{T}}K_{0}B$; for such $B$ (not unique), we further determine which minimizes $K_{2}-B^{\mathrm{T}}K_{1}B$; for such $B$ (still not unique), we finally determine which minimizes $K_{3}-B^{\mathrm{T}}K_{2}B$. This time we can solve $B$ uniquely up to a $\pm$ sign, which can be determined by $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$.
We will derive QWENDY by solving the second problem, and then prove that it also solves the first problem.
3.2 Algorithm details
In this section, given general covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ that may not satisfy Eqs. 13 for any $B_{0}$, we introduce a procedure to calculate $B$ that approximately solves Eqs. 13. Under some mild conditions (some matrices are invertible and have distinct eigenvalues), $B$ can be uniquely determined up to a $\pm$ sign. Then we use $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ to distinguish between $B$ and $-B$. The whole procedure is named as QWENDY method.
Step (1): For any $B$, we want to minimize
$||K_{1}-B^{\mathrm{T}}K_{0}B||_{\mathrm{F}}^{2}.$ (7)
Assume $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ are invertible. Consider Cholesky decompositions
$K_{1}=L_{1}L_{1}^{\mathrm{T}},\ K_{0}=L_{0}L_{0}^{\mathrm{T}},$
where $L_{1}$ and $L_{0}$ are lower-triangular and invertible. Define
$O=L_{1}^{-1}B^{\mathrm{T}}L_{0},$
then the target Eq. 7 becomes
$||L_{1}L_{1}^{\mathrm{T}}-L_{1}OO^{\mathrm{T}}L_{1}^{\mathrm{T}}||_{\mathrm{F}}^{2}.$
Therefore, Eq. 7 is minimized to 00 if and only if $O$ is orthonormal: $OO^{\mathrm{T}}=I$.
Since
$B=L_{0}^{-\mathrm{T}}O^{\mathrm{T}}L_{1}^{\mathrm{T}},$
we use $K_{0}$ and $K_{1}$ to restrict B to a space with the same dimension as the set of all orthonormal matrices.
Step (2): For such $B$ that minimizes Eq. 7, we want to find $B$ that makes $B^{\mathrm{T}}K_{1}B$ close to $K_{2}$. Here we do not minimize
$||K_{2}-B^{\mathrm{T}}K_{1}B||_{\mathrm{F}}^{2}$
as it is difficult. Instead, we want to minimize
$||L_{1}^{-1}(K_{2}-B^{\mathrm{T}}K_{1}B)L_{1}^{-\mathrm{T}}||_{\mathrm{F}}^{2}$ (8)
among $B$ that minimizes Eq. 7.
Assume that $L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}$ does not have repeated eigenvalues. Consider the eigenvalue decomposition decomposition
$L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}=P_{1}D_{1}P_{1}^{\mathrm{T}},$
where $P_{1}$ is orthonormal, and $D_{1}$ is diagonal with strictly increasing positive diagonal elements (eigenvalues), since $L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}$ is positive definite and symmetric. Similarly, assume that $L_{1}^{-1}K_{2}L_{1}^{-\mathrm{T}}$ does not have repeated eigenvalues, and we have
$L_{1}^{-1}K_{2}L_{1}^{-\mathrm{T}}=P_{2}D_{2}P_{2}^{\mathrm{T}}$
with orthonormal $P_{2}$ and diagonal $D_{2}$ with strictly increasing positive diagonal elements (eigenvalues). Now the target Eq. 8 equals
$||P_{2}D_{2}P_{2}^{\mathrm{T}}-OP_{1}D_{1}P_{1}^{\mathrm{T}}O^{\mathrm{T}}||^{2}_{\mathrm{F}}=||D_{2}-P_{2}^{\mathrm{T}}OP_{1}D_{1}P_{1}^{\mathrm{T}}O^{\mathrm{T}}P_{2}||^{2}_{\mathrm{F}},$
since $P_{2}$ is orthonormal and does not affect Frobenius norm. Define
$W=P_{2}^{\mathrm{T}}OP_{1},$
which is orthonormal. Then
$O=P_{2}WP_{1}^{\mathrm{T}},$
and Eq. 8 equals
$||D_{2}-WD_{1}W^{\mathrm{T}}||^{2}_{\mathrm{F}}.$ (9)
We use the following lemma to handle Eq. 9:
Lemma 1.
For diagonal $D_{1},D_{2}$ with strictly increasing diagonal elements and any orthonormal $W$, Eq. 9 is minimized when $W$ is diagonal, and the diagonal elements are ±$\pm 1$. (There are $2^{n}$ possibilities for such $W$.)
Proof.
Denote the diagonal elements of $D_{1}$ as $\mathbf{d}_{1}=[d_{1,1},\ldots,d_{1,n}]$ with $d_{1,1}<d_{1,2}<\cdots,<d_{1,n-1}<d_{1,n}$, and similarly $\mathbf{d}_{2}=[d_{2,1},\ldots,d_{2,n}]$ with $d_{2,1}<d_{2,2}<\cdots,<d_{2,n-1}<d_{2,n}$ for $D_{2}$. We need to minimize the following norm with an orthonormal $W$.
$\begin{split}&||D_{2}-WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}=||D_{2}||_{\mathrm{F}}^{2}+||WD_{1}W^{\mathrm{T}}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}[D_{2}\otimes(WD_{1}W^{\mathrm{T}})]_{ii}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\sum_{i=1}^{n}d_{2,i}\sum_{j=1}^{n}W_{ij}^{2}d_{1,j}\\ =&||D_{2}||_{\mathrm{F}}^{2}+||D_{1}||_{\mathrm{F}}^{2}-2\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}},\end{split}$
where $\otimes$ is the element-wise product. Thus we just need to maximize $\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}}$. Notice that $W\otimes W$ is doubly-stochastic, meaning that it is non-negative, and each row or column has sum $1$. By Birkhoff–von Neumann Theorem [37], $W\otimes W$ can be decomposed to
$W\otimes W=\sum_{i=1}^{k}c_{i}Q_{i},$
where $Q_{i}$ is a permutation matrix, $c_{i}>0$, and $\sum_{i=1}^{k}c_{i}=1$.
Due to the rearrangement inequality [38], for each permutation matrix $Q_{i}$,
$\mathbf{d}_{1}Q_{i}\mathbf{d}_{2}^{\mathrm{T}}\leq\mathbf{d}_{1}I\mathbf{d}_{2}^{\mathrm{T}}=\mathbf{d}_{1}\mathbf{d}_{2}^{\mathrm{T}},$
where the equality holds if and only if $Q_{i}=I$.
Therefore,
$\mathbf{d}_{1}(W\otimes W)\mathbf{d}_{2}^{\mathrm{T}}=\sum_{i=1}^{k}c_{i}\mathbf{d}_{1}Q_{i}\mathbf{d}_{2}^{\mathrm{T}}\leq\sum_{i=1}^{k}c_{i}\mathbf{d}_{1}\mathbf{d}_{2}^{\mathrm{T}}=\mathbf{d}_{1}\mathbf{d}_{2}^{\mathrm{T}},$
where the equality holds if and only if
$W\otimes W=I,$
meaning that $W$ is diagonal, and diagonal elements are ±$\pm 1$. ∎
We now have
$B=L_{0}^{-\mathrm{T}}P_{1}WP_{2}^{\mathrm{T}}L_{1}^{\mathrm{T}},$
meaning that given $K_{0},K_{1},K_{2}$, we can restrict $B$ to $2^{n}$ possibilities.
Step (3): For such $B$ that minimizes Eq. 8 among $B$ that minimizes Eq. 7, we want to find $B$ that makes $B^{\mathrm{T}}K_{2}B$ close to $K_{3}$. Here we do not minimize
$||K_{3}-B^{\mathrm{T}}K_{2}B||_{\mathrm{F}}^{2}$
as it is difficult. Instead, we want to minimize
$||L_{1}^{-1}(K_{3}-B^{\mathrm{T}}K_{2}B)L_{1}^{-\mathrm{T}}||_{\mathrm{F}}^{2}.$ (10)
Define
$G=P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2},$
and
$H=P_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1}.$
Eq. 10 equals
$\begin{split}&||P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2}-WP_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1}W||_{\mathrm{F}}^{2}\\ =&||G-WHW||_{\mathrm{F}}^{2}=||G||_{\mathrm{F}}^{2}+||WHW||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}\\ =&||G||_{\mathrm{F}}^{2}+||H||_{\mathrm{F}}^{2}-2||G\otimes(WHW)||_{\mathrm{F}}^{2}.\end{split}$ (11)
We want to find diagonal $W$ with ±$\pm 1$ that minimizes Eq. 10, which is equivalent to maximizing $||G\otimes(WHW)||_{\mathrm{F}}^{2}$. Define $C=G\otimes H$, which is still positive definite and symmetric by Schur product theorem [39]. Assume that $C$ does not have repeated eigenvalues. Denote the diagonal elements of $W$ by $\mathbf{w}=[w_{1},\ldots,w_{n}]$. Then we need to maximize
$||G\otimes(WHW)||_{\mathrm{F}}^{2}=\mathbf{w}C\mathbf{w}^{\mathrm{T}}.$
Now we relax this problem from $\mathbf{w}$ with $w_{i}=\pm 1$ to general $\mathbf{v}$ with $||\mathbf{v}||_{2}^{2}=n$:
$\max_{||\mathbf{v}||_{2}^{2}=n}\mathbf{v}C\mathbf{v}^{\mathrm{T}}.$
The unique solution (up to a $\pm$ sign) is the eigenvector that corresponds to the largest eigenvalue of $C$.
After obtaining $\mathbf{v}$, we just need to project it to $[w_{1},\ldots,w_{n}]$ with $w_{i}=\pm 1$ by taking the sign of each term: $w_{i}=\mathrm{sign}(v_{i})$. Then construct $W$ by putting $[w_{1},\ldots,w_{n}]$ on the diagonal.
This relaxation does not guarantee obtaining the optimal solution for Eq. 11, but we find that it produces the correct answer in almost all simulations. Alternatively, we can directly determine the optimal $W$ for Eq. 11 by brute-force search, as there are finitely many ($2^{n}$) possibilities of $W$.
With
$B=L_{0}^{-\mathrm{T}}P_{1}WP_{2}^{\mathrm{T}}L_{1}^{\mathrm{T}},$
given $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$, we can uniquely determine $B$ up to a $\pm$ sign. Since $B^{\mathrm{T}}KB=(-B)^{\mathrm{T}}K(-B)$, more $K_{i}$ cannot provide more information.
Step (4): For $B$ and $-B$ from Step (3), we determine which satisfies Eqs. 46 better. Notice that Eqs. 46 share the same unknown $\mathbf{c}$. Define
$\mathbf{c}_{0}=(\mathbf{x}_{1}-\mathbf{x}_{0}B)/t,\ \mathbf{c}_{1}=(\mathbf{x}_{2}-\mathbf{x}_{1}B)/t,\ \mathbf{c}_{2}=(\mathbf{x}_{3}-\mathbf{x}_{2}B)/t,\ \bar{\mathbf{c}}=(\mathbf{c}_{0}+\mathbf{c}_{1}+\mathbf{c}_{2})/3.$
Then the error for fitting Eqs. 46 is
$||\mathbf{c}_{0}-\bar{\mathbf{c}}||_{2}^{2}+||\mathbf{c}_{1}-\bar{\mathbf{c}}||_{2}^{2}+||\mathbf{c}_{2}-\bar{\mathbf{c}}||_{2}^{2}=||\mathbf{c}_{0}||_{2}^{2}+||\mathbf{c}_{1}||_{2}^{2}+||\mathbf{c}_{2}||_{2}^{2}-3||\bar{\mathbf{c}}||_{2}^{2}.$
We just need to compare the errors of $B$ and $-B$ and choose the smaller one as the output.
With expression data from four time points, we uniquely determine $B$. See Algorithm 1 for the workflow of QWENDY method.
  1. 1.
    Input expression levels of $n$ genes over $m$ cells at time points 00, $t$, $2t$, $3t$
  2. 2.
    Calculate covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$, and mean levels $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$
  3. 3.
    Calculate Cholesky decomposition
    $K_{1}=L_{1}L_{1}^{\mathrm{T}},\ K_{0}=L_{0}L_{0}^{\mathrm{T}}$
    Calculate eigenvalue decomposition
    $L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}=P_{1}D_{1}P_{1}^{\mathrm{T}},\ L_{1}^{-1}K_{2}L_{1}^{-\mathrm{T}}=P_{2}D_{2}P_{2}^{\mathrm{T}}$
    Calculate
    $C=(P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2})\otimes(P_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1})$
  4. 4.
    Calculate $\mathbf{v}$, the eigenvector that corresponds to the largest eigenvalue of $C$, and the projection $\mathbf{w}$ with $w_{i}=\mathrm{sign}(v_{i})$
    Construct $W$ with $\mathbf{w}$ on diagonal
  5. 5.
    Calculate $B=L_{0}^{-\mathrm{T}}P_{1}WP_{2}^{\mathrm{T}}L_{1}^{\mathrm{T}}$
    Compare total squared errors of $B$ and $-B$ for Eqs. 46
  6. 6.
    Output $B$ or $-B$, the one that corresponds to the smaller error, and the GRN $A=(B-I)/t$
Algorithm 1 Workflow of QWENDY method.
3.3 Correctness of QWENDY method
Theorem 1.
If covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and mean levels $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ satisfy Eqs. 16 for some $B_{0}$, then QWENDY will output $B_{0}$.
Proof.
If $K_{1}=B_{0}^{\mathrm{T}}K_{0}B_{0}$, then
$(L_{1}^{-1}B_{0}^{\mathrm{T}}L_{0})(L_{0}^{\mathrm{T}}B_{0}L_{1}^{-\mathrm{T}})=I,$
and $L_{1}^{-1}B_{0}^{\mathrm{T}}L_{0}=O_{0}$ for some orthonormal $O_{0}$. Thus
$B_{0}=L_{0}^{-\mathrm{T}}O_{0}^{\mathrm{T}}L_{1}^{\mathrm{T}},$
which is among the calculated $B$ in Step (1).
If $K_{2}=B_{0}^{\mathrm{T}}K_{1}B_{0}$, then
$D_{2}=P_{2}^{\mathrm{T}}O_{0}P_{1}D_{1}P_{1}^{\mathrm{T}}O_{0}^{\mathrm{T}}P_{2}.$
Since $P_{2}^{\mathrm{T}}O_{0}P_{1}$ is orthonormal, $D_{1}$ and $D_{2}$ are similar and have the same eigenvalues. Therefore, $D_{1}$ and $D_{2}$ as diagonal matrices with increasing diagonal values are equal. Assume that the diagonal elements $d_{1},\ldots,d_{n}$ of $D_{1}$ are distinct. Define $W_{0}=P_{2}^{\mathrm{T}}O_{0}P_{1}$, then
$D_{1}W_{0}=W_{0}D_{1},$
and $W_{0}[i,j]d_{i}=W_{0}[i,j]d_{j}$. For $i\neq j$, we have $d_{i}\neq d_{j}$, meaning that $W_{0}[i,j]=0$. Now $W_{0}$ is diagonal and orthonormal, implying that its diagonal elements are $1$ or $-1$, and
$B_{0}=L_{0}^{-\mathrm{T}}P_{1}W_{0}P_{2}^{\mathrm{T}}L_{1}^{\mathrm{T}}$
is among the calculated $B$ in Step (2).
If $K_{3}=B_{0}^{\mathrm{T}}K_{2}B_{0}$, then
$P_{2}^{\mathrm{T}}L_{1}^{-1}K_{3}L_{1}^{-\mathrm{T}}P_{2}=W_{0}P_{1}^{\mathrm{T}}L_{0}^{-1}K_{2}L_{0}^{-\mathrm{T}}P_{1}W_{0}.$
Define $\mathbf{w}_{0}=[w_{1},\ldots,w_{n}]$ to be the diagonal elements of $W_{0}$. Then from Step (3), $W_{0}$ minimizes Eq. 10, and $\mathbf{w}_{0}$ is the solution to
$\max_{||\mathbf{v}||_{2}^{2}=n}\mathbf{v}C\mathbf{v}^{\mathrm{T}},$
namely the eigenvector of the largest eigenvalue. Since $w_{i}=\pm 1$, the projection $w_{i}=\mathrm{sign}(v_{i})$ has no effect, and
$B_{0}=L_{0}^{-\mathrm{T}}P_{1}W_{0}P_{2}^{\mathrm{T}}L_{1}^{\mathrm{T}}$
is the unique $B$ (up to a $\pm$ sign) calculated in Step (3).
From Eqs. 46,
$\mathbf{x}_{1}-\mathbf{x}_{0}B_{0}=\mathbf{x}_{2}-\mathbf{x}_{1}B_{0}.$
Assume $\mathbf{x}_{1}\neq\mathbf{x}_{2}$, then
$\mathbf{x}_{1}-\mathbf{x}_{0}(-B_{0})\neq\mathbf{x}_{2}-\mathbf{x}_{1}(-B_{0}).$
Therefore, $-B_{0}$ does not satisfy Eqs. 46, and Step (4) chooses the correct $B_{0}$ from the two possibilities in Step (3). ∎
4 Enhanced QWENDY method
QWENDY method is derived from Eqs. 16, especially Eqs. 13, which are a linear approximation of the actual nonlinear gene expression dynamics. Therefore, real $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ might not fit with Eqs. 13, and it might not be feasible to apply QWENDY directly to real $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$. This problem already exists for WENDY method, where the input matrices $K_{0}$, $K_{1}$ might not satisfy
$K_{1}=B^{\mathrm{T}}K_{0}B.$
For this problem, TRENDY method [10] proposes that we can construct $K_{1}^{*}=B^{\mathrm{T}}K_{0}B$, and train a model with input $K_{1}$ and target $K_{1}^{*}$, so that the output $K_{1}^{\prime}$ is close to $K_{1}^{*}$. Then we can apply WENDY to $K_{0}$, $K_{1}^{\prime}$ to obtain a more accurate GRN $A_{1}$. After that, we can train another model with input $K_{0}$, $K_{1}$, $A_{1}$, and the target is the true GRN $A_{\text{true}}$. Then the final output $A_{2}$ is more similar to the true GRN than $A_{1}$.
Inspired by TRENDY, we propose a similar solution to enhance QWENDY. See Algorithms 2,3 for details. Since $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ are only used to distinguish between $B$ and $-B$, it is not necessary to train another model for them.
  1. 1.
    Repeat generating random GRN $A_{\mathrm{true}}$ and corresponding gene expression data at time 00, $t$, $2t$, $3t$ from a gene expression simulator
  2. 2.
    Calculate covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and mean levels $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$, and then calculate
    $K_{0}^{*}=K_{0},\ K_{1}^{*}=B^{\mathrm{T}}K_{0}^{*}B,\ K_{2}^{*}=B^{\mathrm{T}}K_{1}^{*}B,\ K_{3}^{*}=B^{\mathrm{T}}K_{2}^{*}B,$
    where $B=I+tA_{\mathrm{true}}$
  3. 3.
    Train matrix-learning model 1 with inputs $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and target $K_{0}^{*}$, $K_{1}^{*}$, $K_{2}^{*}$, $K_{3}^{*}$
    Call trained matrix-learning model 1 to calculate $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$ from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$
  4. 4.
    Call QWENDY to calculate $A_{1}$ from $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$, $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$
  5. 5.
    Train matrix-learning model 2 with input $A_{1}$, $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and target $A_{\mathrm{true}}$
Algorithm 2 Training workflow of LEQWENDY/TEQWENDY method. For those two matrix-learning models, LEQWENDY adopts LE structure, and TEQWENDY adopts TE structure.
  1. 1.
    Input: gene expression data at four equally spaced time points
  2. 2.
    Calculate covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ and mean levels $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$
  3. 3.
    Call trained matrix-learning model 1 to calculate $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$ from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$
  4. 4.
    Call QWENDY to calculate $A_{1}$ from $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$, $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$
  5. 5.
    Call trained matrix-learning model 2 to calculate $A_{2}$ from $A_{1}$, $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$
  6. 6.
    Output: inferred GRN $A_{2}$
Algorithm 3 Testing workflow of LEQWENDY/TEQWENDY method. For those two matrix-learning models, LEQWENDY adopts LE structure, and TEQWENDY adopts TE structure.
For those two matrix-learning models in Algorithms 2,3, we can adopt the approach of TRENDY to construct a transformer structure with three sections: (1) Pre-process the inputs; (2) Use transformer encoder layers to learn high-dimensional representations of the inputs; (3) Construct outputs from the high-dimensional representations.
For training the transformer encoder layers in this structure, we present two approaches. One approach is to train new transformer encoder layers from scratch, the same as TRENDY. This structure is named “TE”, meaning “transformer-enhanced”.
The other approach is to integrate an LLM and fine-tune it. In general, an LLM has three sections: (1) Convert text to vector representations; (2) Use transformer (encoder layers, decoder layers, or both) to learn contextual relationships; (3) Convert model outputs back to natural language [40]. We can choose an LLM with pre-trained transformer encoder layers in the second section, and use them to replace the transformer encoder layers in the TE structure. In our practice, we use the encoder layers of the RoBERTa-large model [41], which have about 300 million parameters.
The pre-trained parameters will be kept frozen while we adopt a parameter-efficient fine-tuning method: LoRA [42]. Lower-rank matrices will be trained and added to the frozen encoding layers to obtain a new encoder in a cost-efficient way. The new structure with RoBERTa is named “LE”, representing “LLM-enhanced”.
We name Algorithms 2,3 with TE structure as TEQWENDY, and Algorithms 2,3 with LE structure as LEQWENDY. See Appendix for details of TE and LE structures. TEQWENDY has 4.7 million trainable parameters. LEQWENDY has 4.6 million trainable parameters, with 300 million non-trainable (frozen) parameters.
TEQWENDY and LEQWENDY are trained on synthetic data generated by [43, 8, 9, 10]
$\mathrm{d}X_{j}(t)=V\left\{\beta\prod_{i=1}^{n}\left[1+(A_{\mathrm{true}})_{i,j}\frac{X_{i}(t)}{X_{i}(t)+1}\right]-\theta X_{j}(t)\right\}\mathrm{d}t+\sigma X_{j}(t)\mathrm{d}W_{j}(t),$ (12)
where $X_{i}(t)$ is the level of gene $i$ at time $t$, $W_{j}(t)$ is a standard Brownian motion, and $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$ There are $10^{5}$ training samples, each with the expression levels of 10 genes for 100 cells, measured at four time points.
QWENDY requires that $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ are symmetric and positive definite. Although the input $K_{i}$ and the target $K_{i}^{*}$ are naturally symmetric and positive definite, the learned output $K_{i}^{\prime}$ in Step (3) of Algorithms 2,3 might not be positive definite, or even symmetric. Therefore, in the implementation of QWENDY, we add two extra steps to adjust the input covariance matrices:
1. If $K_{i}$ is asymmetric, replace $K_{i}$ by $(K_{i}+K_{i}^{\mathrm{T}})/2$.
2. If $K_{i}$ is not positive definite, in the eigenvalue decomposition $K_{i}=O\Lambda O^{\mathrm{T}}$, where $O$ is orthonormal, and $\Lambda$ is diagonal, replace negative values of $\Lambda$ by small positive values to obtain $\Lambda^{\prime}$, and replace $K_{i}$ by $O\Lambda^{\prime}O^{\mathrm{T}}$.
5 Performance of different methods
The same as in the WENDY paper and the TRENDY paper [9, 10], we test different methods on two synthetic data sets and two experimental data sets. SINC data set [10] and DREAM4 data set [44] are generated by simulating stochastic differential equation systems. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [45]. hESC data set is from human embryonic stem cell-derived progenitor cells[46]. THP-1 and hESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
For SINC data set, we use data from all four time points. For DREAM4 data set, we use data from any four consecutive time points and take average. For THP-1 data set and hESC data set, we choose any four time points with equal difference. This is due to the limitation that QWENDY (and thus TEQWENDY and LEQWENDY) is derived for four evenly spaced time points.
We test the performance of QWENDY, TEQWENDY, and LEQWENDY methods on these four data sets. Since TRENDY ranks the first among 16 methods on these four data sets [10], we also present the performance of TRENDY. For each method, we compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores [9]. These two scores, both between 00 and $1$, evaluate the level of matching, where $1$ means perfect match, and 00 means perfect mismatch.
See Table 1 for the scores. We can see that TEQWENDY is slightly better than TRENDY, meaning that TEQWENDY has the best performance in all 19 methods (16 methods tested in TRENDY paper, QWENDY, TEQWENDY, and LEQWENDY). Besides, QWENDY is worse than TRENDY, and LEQWENDY is even worse than QWENDY, showing that LLMs trained with natural language inputs might not directly help with the numerical task in GRN inference.
Since TEQWENDY and TRENDY are trained on synthetic data, a natural concern is whether the training makes these methods overfitting to certain types of synthetic data, but less suitable for experimental data. Therefore, we also compare the performance of different methods only for two experimental data sets. For all 16 methods tested in TRENDY paper, bWENDY ranks the first on THP-1 and hESC data sets, with total score 1.5783 [10]. In comparison, QWENDY has total score 1.6272, making it the best in all 19 methods on two experimental data sets.
In sum, TEQWENDY method has the best overall performance, while QWENDY ranks the first on experimental data sets. LEQWENDY is worse than QWENDY in most cases, showing the failure of training with an LLM.
QWENDY
LEQ-
WENDY
TEQ-
WENDY
TRENDY
SINC AUROC 0.5314 0.4972 0.9587 0.8941
AUPRC 0.5720 0.5185 0.9142 0.8314
DREAM4 AUROC 0.4987 0.5164 0.5372 0.5341
AUPRC 0.1844 0.1823 0.2203 0.2177
THP-1 AUROC 0.5524 0.5543 0.5415 0.5557
AUPRC 0.4294 0.3632 0.3801 0.3669
hESC AUROC 0.6019 0.5905 0.4815 0.5311
AUPRC 0.0435 0.0367 0.0317 0.0376
Total 3.4137 3.2591 4.0652 3.9686
Table 1: AUROC and AUPRC scores of QWENDY, LEQWENDY, TEQWENDY, TRENDY methods on four data sets
6 Discussion
In this paper, we present QWENDY, a GRN inference method that requires single-cell gene expression data measured at four time points. Then we train the transformer-based variant TEQWENDY from scratch and fine-tune the LLM-based variant LEQWENDY based on pre-trained weights. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets. Notice that each experimental data set only has one group of data, meaning that the performance of each method has a large uncertainty level.
The performance of LEQWENDY is not satisfactory, meaning that the pre-trained LLMs, even after task-specific fine-tuning, do not fit with this task. We need to explore other approaches of applying LLMs in GRN inference, and utilize the fact that LLMs are trained with natural language material.
QWENDY method and its variants require measurements at four evenly spaced time points $t_{0}$, $t_{1}$, $t_{2}$, $t_{3}$, meaning that $t_{1}-t_{0}=t_{2}-t_{1}=t_{3}-t_{2}$. Otherwise, the dynamics of covariance matrices is much more complicated, and we do not have an explicit solution. One possible strategy is to use the approximation
$I+(t_{2}-t_{1})A\approx\frac{t_{2}-t_{1}}{t_{1}-t_{0}}\left[I+(t_{1}-t_{0})A\right],$
so that different $B$ in Eqs. 1-3 only differ by a constant factor. This means that we can relocate this factor to $K_{i}$, and $B$ is the same for all equations.
Since QWENDY and its variants need four time points, the duration of the whole experiment might be too long, so that the dynamics of gene expression under regulation might have changed during this time. QWENDY is based on a time-homogeneous model, which might fail in this situation.
Training TEQWENDY and LEQWENDY on data generated by Eq. 12 does not necessarily increase their performance on experimental data sets. This means that Eq. 12 might not faithfully reflect the gene expression dynamics. To better integrate deep learning techniques, we need better gene expression data generators.
Although QWENDY and its variants can fully determine the GRN, the diagonal elements that represent the regulation of one gene to itself (autoregulation) can be mixed with natural mRNA synthesis and degradation [47]. Therefore, it is not recommended to infer the autoregulation with QWENDY or its variants.
Given new GRN inference methods, we can study how cells maintain homeostasis [48, 49] or be driven away from homeostasis and cause diseases [50, 51].
Acknowledgments
Y.W. would like to thank Dr. Mingda Zhang for a helpful discussion.
Data and code availability
Main function of QWENDY, TEQWENDY, and LEQWENDY methods, including a tutorial and the training files, along with other data and code files used in this paper, can be found in
https://github.com/YueWangMathbio/QWENDY
Appendix: details of TEQWENDY and LEQWENDY
For all models in TEQWENDY and LEQWENDY, the loss function is mean squared error; the optimizer is Adam with learning rate $0.001$; the number of training epochs is $100$. After each epoch, we evaluate the model performance on a validation set with 1000 samples. The training stops early and rolls back to the best status if there is no improvement for consecutive $10$ epochs.
See Algorithm 4 for the structure of the first half of TEQWENDY. The inputs are four covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$. The targets are four revised covariance matrices $K_{0}^{*}$, $K_{1}^{*}$, $K_{2}^{*}$, $K_{3}^{*}$. Notice that the first input matrix ($K_{0}$) is not processed through these layers, since the target $K_{0}^{*}$ equals $K_{0}$. Besides, $K_{1}$, $K_{2}$, $K_{3}$ are processed separately.
  1. 1.
    Input: four covariance matrices ($4$ groups of $n\times n$)
  2. 2.
    Linear embedding layer with dimension $1$ to $d=64$ ($4$ groups of $n\times n\times d$)
  3. 3.
    ReLU activation function ($4$ groups of $n\times n\times d$)
  4. 4.
    Linear embedding layer with dimension $d$ to $d$ ($4$ groups of $n\times n\times d$)
  5. 5.
    2-D positional encoding layer ($4$ groups of $n\times n\times d$)
  6. 6.
    Flattening and concatenation ($4$ groups of $(n^{2})\times d$)
  7. 7.
    $7$ layers of transformer encoder with $4$ heads, dimension $d$, feedforward dimension $4d$, dropout rate $0.1$ ($4$ groups of $(n^{2})\times d$)
  8. 8.
    Linear embedding layer with dimension $d$ to $d$ ($4$ groups of $(n^{2})\times d$)
  9. 9.
    LeakyReLU activation function with $\alpha=0.1$ ($4$ groups of $n\times n\times d$)
  10. 10.
    Linear embedding layer with dimension $d$ to $1$ ($4$ groups of $(n^{2})\times 1$)
  11. 11.
    Output: reshaping into four matrices ($4$ groups of $n\times n$)
Algorithm 4 Structure of TEQWENDY method, first half. The shape of data after each layer is in the brackets.
The 2-D positional encoding layer incorporates spatial information of matrix to the input. It generates an $n\times n\times d$ array PE and adds it to the embedded input: For $x$ and $y$ in $1,2,\ldots,n$ and $j$ in $41,\ldots,d/$ 4,
$\begin{split}&\text{PE}[x,y,2j-1]=\cos[(x-1)\times 10^{-16(j-1)/d}],\\ &\text{PE}[x,y,2j]=\sin[(x-1)\times 10^{-16(j-1)/d}],\\ &\text{PE}[x,y,2j-1+d/2]=\cos[(y-1)\times 10^{-16(j-1)/d}],\\ &\text{PE}[x,y,2j+d/2]=\sin[(y-1)\times 10^{-16(j-1)/d}].\end{split}$
See Algorithm 5 for the structure of the second half of TEQWENDY. After obtaining the outputs $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$ from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ by the first half of TEQWENDY, call the QWENDY method to calculate the inferred GRN $A_{1}$ from $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$. The inputs of the second half of TEQWENDY are four covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$, and the inferred GRN $A_{1}$. The target is the ground truth GRN $A_{\mathrm{true}}$.
  1. 1.
    Input: four covariance matrices and one inferred GRN ($5$ groups of $n\times n$)
  2. 2.
    Linear embedding layer with dimension $1$ to $d=64$ ($5$ groups of $n\times n\times d$)
  3. 3.
    Segment embedding layer ($5$ groups of $n\times n\times d$)
  4. 4.
    2-D positional encoding layer ($5$ groups of $n\times n\times d$)
  5. 5.
    Flattening and concatenation ($(n^{2})\times(5d)$)
  6. 6.
    $3$ layers of transformer encoder with $4$ heads, dimension $5d$, feedforward dimension $20d$, dropout rate $0.1$ ($(n^{2})\times(5d)$)
  7. 7.
    Linear embedding layer with dimension $5d$ to $1$ ($(n^{2})\times 1$)
  8. 8.
    Output: reshaping into one matrix ($n\times n$)
Algorithm 5 Structure of TEQWENDY method, second half. The shape of data after each layer is in the brackets.
The segment embedding layer generates different trainable $d$-dimensional vectors for all five inputs. Then each vector is copied into dimension $n\times n\times d$, and added to the embedded inputs. This layer incorporates the source of inputs. The final input of the transformer encoder layers is a matrix, with shape $k\times D$, where $k$ is the total number of input values, and $D$ is the representation dimension. For each location in $1,2,\ldots,k$, the segment embedding marks which input matrix it is from, and the position encoding marks which position in the matrix it is from. These two layers solve the problem that the inputs are multiple matrices, but the inputs of transformer are representations of a 1-D sequence.
See Algorithm 6 for the structure of the first half of LEQWENDY. The inputs are four covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$. The targets are four revised covariance matrices $K_{0}^{*}$, $K_{1}^{*}$, $K_{2}^{*}$, $K_{3}^{*}$. Notice that the first input matrix ($K_{0}$) is not processed through these layers, since the target $K_{0}^{*}$ equals $K_{0}$. Besides, $K_{1}$, $K_{2}$, $K_{3}$ are processed separately.
  1. 1.
    Input: four covariance matrices ($4$ groups of $n\times n$)
  2. 2.
    Linear embedding layer with dimension $1$ to $d=256$ ($4$ groups of $n\times n\times d$)
  3. 3.
    ReLU activation function ($4$ groups of $n\times n\times d$)
  4. 4.
    Linear embedding layer with dimension $d$ to $d$ ($4$ groups of $n\times n\times d$)
  5. 5.
    Segment embedding layer ($4$ groups of $n\times n\times d$)
  6. 6.
    2-D positional encoding layer ($4$ groups of $n\times n\times d$)
  7. 7.
    Flattening and concatenation ($(n^{2})\times(4d)$)
  8. 8.
    Transformer encoder part of the RoBERTa-large model, frozen: $24$ layers of transformer encoder with $16$ heads, dimension $4d$, feedforward dimension $16d$, dropout rate $0.1$;
[…]
  9. 9.
    Four different linear embedding layers with dimension $4d$ to $2d$ ($4$ groups of $(n^{2})\times(2d)$)
  10. 10.
    LeakyReLU activation function with $\alpha=0.1$ ($4$ groups of $(n^{2})\times(2d)$)
  11. 11.
    Four different linear embedding layers with dimension $2d$ to $1$ ($4$ groups of $(n^{2})\times 1$)
  12. 12.
    Output: reshaping into four matrices ($4$ groups of $n\times n$)
Algorithm 6 Structure of LEQWENDY method, first half. The shape of data after each layer is in the brackets.
[…]
See Algorithm 7 for the structure of the second half of LEQWENDY. After obtaining the outputs $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$ from $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ by the first half of LEQWENDY, call the QWENDY method to calculate the inferred GRN $A_{1}$ from $K_{0}^{\prime}$, $K_{1}^{\prime}$, $K_{2}^{\prime}$, $K_{3}^{\prime}$. The inputs of the second half of LEQWENDY are four covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$, and the inferred GRN $A_{1}$. The target is the ground truth GRN $A_{\mathrm{true}}$. Since the transformer encoder part of the RoBERTa-large model has a fixed dimension 1024, we need to apply $d_{1}=192$ for each covariance matrix input, and $d_{2}=256$ for the GRN input, so that the total dimension is $d=4d_{1}+d_{2}=1024$.
  1. 1.
    Input: four covariance matrices and one inferred GRN ($5$ groups of $n\times n$)
  2. 2.
    Linear embedding layer with dimension $1$ to $d_{1}=192$ or $d_{2}=256$ ($4$ groups of $n\times n\times d_{1}$ and $1$ group of $n\times n\times d_{2}$)
  3. 3.
    Segment embedding layer ($4$ groups of $n\times n\times d_{1}$ and $1$ group of $n\times n\times d_{2}$)
  4. 4.
    2-D positional encoding layer ($4$ groups of $n\times n\times d_{1}$ and $1$ group of $n\times n\times d_{2}$)
  5. 5.
    Flattening and concatenation ($(n^{2})\times(d=4d_{1}+d_{2}=1024)$)
  6. 6.
    Transformer encoder part of the RoBERTa-large model, frozen: $24$ layers of transformer encoder with $16$ heads, dimension $4d$, feedforward dimension $16d$, dropout rate $0.1$;
[…]
  7. 7.
    Linear embedding layer with dimension $d$ to $d$ ($(n^{2})\times d$)
  8. 8.
    LeakyReLU activation function with $\alpha=0.1$ ($(n^{2})\times d$)
  9. 9.
    Linear embedding layer with dimension $d$ to $1$ ($(n^{2})\times 1$)
  10. 10.
    Output: reshaping into one matrix ($n\times n$)
Algorithm 7 Structure of LEQWENDY method, second half. The shape of data after each layer is in the brackets.